### v2

# Driving by Hybrid Navigation: An Online HD-SD Map Association Framework and Benchmark for Autonomous Vehicles

Abstract
Autonomous vehicles rely on global standard-definition (SD) maps for road-level route planning and online local high-definition (HD) maps for lane-level navigation. However, recent work concentrates on construct online HD maps, often overlooking the association of global SD maps with online HD maps for hybrid navigation, making challenges in utilizing online HD maps in the real world. Observing the lack of the capability of autonomous vehicles in navigation, we introduce Online Map Association, the first benchmark for the association of hybrid navigation-oriented online maps, which enhances the planning capabilities of autonomous vehicles. Based on existing datasets, the OMA contains 480k of roads and 260k of lane paths and provides the corresponding metrics to evaluate the performance of the model. Additionally, we propose a novel framework, named Map Association Transformer, as the baseline method, using path-aware attention and spatial attention mechanisms to enable the understanding of geometric and topological correspondences. The code and dataset can be accessed at https://github.com/WallelWan/OMA-MAT.
1 Introduction
As autonomous driving advances rapidly, the researcher’s focus has been on developing precise and reliable navigation systems. The Standard Definition Map (SD Map) and the High Definition Map (HD Map) served as crucial components in navigation systems, each providing different degrees of environmental detail [9]. The SD Map serves as global navigation at the road level, enjoying widespread application due to its frequent updates and economical storage demands [23], despite its meters-level precision and absence of lane specifics, as illustrated in Fig. 1 (a). On the other hand, HD Map provides detailed geometric information and topology at lane-level to facilitate exact positioning and lane-specific route planning [19], but its extensive production costs and slower update cycles restrict its applicability.
Figure 1: The comparison of navigation on (a) SD map , (b) online HD map and (c) SD-HD map association (Ours). (d) The challenge of map association, including GPS shift, missing and noise of centerline. The identical color in both road and centerline indicates a corresponding pair.
Recently, to address the cost problems of global HD maps, the focus has shifted to online HD map construction [17, 14], aimed at constructing localized HD maps surrounding the vehicle using the vehicle’s own sensing system (refer to Fig. 1 (b)). However, because online HD maps prioritize the view of the ego-vehicle, they are unable to provide a complete centerline topology from start to end, which means they lack the capability for navigation [13]. To enable navigation using the online HD map, it is necessary to convert global navigation at the road-level on the SD map into a local lane-level control strategy corresponding to the online HD map [12], which involves creating an association between the road in the SD map and the lane on the online HD map, as illustrated in Fig. 1 (c). The end result is the successful implementation of hybrid navigation that is based on both the SD map and the online HD map.
However, three significant challenges are encountered during map association [19]. Initially, vehicle positioning error or accuracy issues with SD Map can lead to a GPS shift in location between HD Map and SD Map (as depicted in Fig. 1 (d) (1)), causing location-based KNN matching [5] to frequently produce inaccurate results. Secondly, due to obstructions or disturbances, online HD maps face challenges with intricate topology arising from missing (Fig. 1 (d) (2)) or noise (Fig. 1 (d) (3)) of centerline. Hidden Markov Models (HMM) [25] are ineffective here. Third, a significant challenge in this field is the scarcity of large datasets, which has limited deep learning research in this domain.
To address this fundamental challenge, we thoroughly explore three aspects of contribution: (1) dataset construction, (2) evaluation metric, and (3) algorithm development. Using nuScenes, we introduce the first dataset for the online map association (OMA) with an evaluation metric named Association P-R. OMA derived from nuScenes [3] and OpenStreetMap [1] and boasts more than 30k scenarios and more than 480k roads and 260k lane-level paths. Through a combination of automated and manually calibrated annotations, high-precision road-lane matching annotations are achieved. For the evaluation metrics, we introduce a specific metric called Association P-R, which includes three separate metrics: A-P, A-R, and A-F1. This metric is designed for the association of the online map, taking into account both the accuracy and the precision of the topological alignment.
Furthermore, we present the Map Association Transformer (MAT), a transformer-inspired framework designed to form an online map association utilizing dual attention mechanisms. MAT handles vectorized map inputs by leveraging path-aware attention for learn topological representations and spatial attention to aggregate spatial context. With path-aware and spatial attention, the model is capable of simultaneously learning local geometric correspondences along with global topological structures.
In summary, our contributions are three folds:
  • We introduce Online Map Association (OMA), the first benchmark for hybrid navigation-oriented online map association.
  • We introduce Association P-R, a metric for map association that considers the accuracy and precision of topological alignment.
  • We propose a Map Association Transformer (MAT), which utilizes path-aware attention and spatial attention mechanisms to enable understanding of geometric and topological correspondences.
2 Related Work
Path Planning. Planning the path for autonomous driving requires coordinated global and local strategies. Global planning identifies optimal routes using graph-based methods such as Dijkstra [8] and A* [10] on SD maps, while local trajectory prediction generates detailed paths through optimal control algorithms [32, 33]. Current datasets [4, 30] focus on local forecasting based on HD maps but lack SD-HD integration. We extend nuScenes [4] with OpenStreetMap SD links and annotate lane-to-link connections, creating the first HD2SD binding dataset.
Online HD Map Construction. The construction of HD maps is a popular topic in autonomous driving and is crucial for subsequent tasks [35, 29, 34]. HDMapNet [14] pioneered BEV-based map generation through sensor fusion, while LSS [26] introduced depth-aware BEV transformation. VectorMapNet [20] enabled end-to-end vector prediction, and the MapTR series [16, 17] introduced hierarchical query embeddings for instance-level construction. We adapt MapTRv2 for online HD construction using our dataset’s annotations and demonstrate its compatibility with SD-HD association.
Map Association. Associating HD lane centerlines to SD road links provides a global context for trajectory prediction. Conventional methods like HMM [25] struggle with lane-level accuracy, while closed PDM [6] lacks explicit map association. We propose a contrastive learning framework that extracts semantic features from HD maps and explicitly associates them with SD links, establishing a robust baseline for cross-map alignment.
3 Task Definition
Figure 2: (a) The schema of SD map input:Road $\mathcal{R}$. (b) The schema of online HD map input: Centerline $\mathcal{L}$ and boundary $B$. (c) The objective in our task: Mapping function $f$.
In autonomous driving systems, the precise alignment of standard definition (SD) maps with online high definition (HD) maps is critical for lane-level navigation [9, 19]. This task enables planning modules to execute accurate maneuvers by combining real-time HD map observations (e.g., dynamic centerline configurations) with static SD map topologies. We formalize this problem by learning a function that aligns each HD map centerline with its corresponding road in the SD map.
SD Map. As shown in Fig.2 (a), SD maps represent road networks that use roads as primary primitives. Formally, a SD map is defined as a graph $\mathcal{G}_{\mathcal{R}}=(\mathcal{R},\mathcal{E}_{\mathcal{R}})$, where =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{m}\}$ denotes a set of roads and $\mathcal{E}_{\mathcal{R}}$ encodes their topological connectivity. Each road $r_{j}$ is parameterized by a sequence of directed vectors:
$r_{j}=\left(\overrightarrow{q_{j1}q_{j2}},\overrightarrow{q_{j2}q_{j3}},\ldots,\overrightarrow{q_{jk-1}q_{jk}}\right),\quad q_{jk}\in\mathbb{R}^{2},$ (1)
where consecutive points define road segments through uniform spatial sampling.
HD Map. As shown in Fig.2 (b), Online HD maps provide details of the lane level, primarily represented as a center line network. We model a HD map as a graph $\mathcal{G}_{\mathcal{L}}=(\mathcal{L},\mathcal{E}_{\mathcal{L}})$, where =$\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$ is a set of centerlines, each sampled at uniform intervals:
$l_{i}=\overrightarrow{p_{i}^{1}p_{i}^{2}},\quad p_{i}^{1},p_{i}^{2}\in\mathbb{R}^{2}.$ (2)
$\mathcal{E}_{\mathcal{L}}$ captures topological relations between the adjacent centerlines. In addition, we include road boundary vectors =$\mathcal{B}=\{b_{1},b_{2},\ldots,b_{m_{b}}\}$, which reflect the extent and shape of the actual road. Each boundary $b_{j}$ is specified as:
$b_{j}=\left(\overrightarrow{h_{j1}h_{j2}},\overrightarrow{h_{j2}h_{j3}},\ldots,\overrightarrow{h_{jk-1}h_{jk}}\right),\quad h_{jk}\in\mathbb{R}^{2}.$ (3)
Objective. As shown in Fig.2 (c), given $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{L}}$, the task is to learn a mapping function $f:\mathcal{L}\rightarrow\mathcal{R}$ that assigns each centerline $l\in\mathcal{L}$ to its corresponding road $r_{l}\in\mathcal{R}$. The function satisfies two key constraints:
  1. 1.
    Uniqueness: Each centerline $l$ maps to exactly one ground truth road $r_{l}$;
  2. 2.
    Multiplicity: A single road $r\in\mathcal{R}$ may be associated with multiple centerlines $l_{1},l_{2},\ldots\in\mathcal{L}$.
This formulation casts the alignment task as a many-to-one classification problem, where the number of classes equals the number of $|\mathcal{R}|$, and each centerline acts as an input sample. The goal is to maximize classification accuracy while preserving topological consistency between HD and SD maps.
4 Dataset and Metric
In this section, we provide a summary of the Online Map Association (OMA) dataset. The Online Map Association (OMA) dataset is built on nuScenes [4] and OpenStreetMap [1] (CC BY-NC-SA 4.0), containing 30K+ HD-SD map pairs with high-quality annotations generated via automated labeling, manual refinement, and multi-stage validation. Dataset statistics are summarized in supplemental material.
Figure 3: (a) The visualization of SD map and HD map with association annotations of Boston in nuScenes. The same color implies an associative pair. (b) The visualization of OMA. (c) The visualization of OMA-GT.
4.1 Dataset Construction
Raw Data and Annotation. The source of raw HD maps is nuScenes [4], which includes locations in Boston and Singapore, featuring centerline geometries scanned with LiDAR. For these areas, SD maps were obtained from OpenStreetMap (OSM). Initially, GPS coordinates are used to align HD and SD maps, after which manual adjustments are made to address any remaining misalignments due to GPS inaccuracies or differences in map projections. The visualization is shown in Fig. 3 (a), which shows the global SD and HD map of Boston after map association.
An automated labeling pipeline using a Map Association Transformer (MAT, Section 5) trained on proprietary Chinese HD-SD maps generates initial annotations on nuScenes, followed by topological post-processing. Professional annotators refine these drafts through geometry and topology adjustments under domain-specific rules. The result is high-quality annotations that meet real-world deployment needs.
OMA. To evaluate robustness to predicted HD maps, the online HD map in OMA is generated by MapTRv2 [17], as shown in Fig 3 (b). The model is trained on nuScenes using synchronized LiDAR and camera inputs with offical configuration. For each sample:
  • SD Map: An SD map cropping measure $150\,m\times 150\,m$ centered around the ego vehicle preserves the adjacent topological context.
  • HD Map: An HD map cropping measure $30\,m\times 60\,m$ centered around the ego vehicle, as referenced by [17, 21, 15], is used as input for the online HD map.
OMA-GT. We propose OMA-GT to further assess the robustness of association algorithms against map dependencies. Unlike OMA, OMA-GT is designed by emulating online perception through the use of localized ground truth HD maps surrounding the ego vehicle, as depicted in Fig 3 (c).
Figure 4: Example of TP, FP and FN for evaluate Association Precision-Recall.
4.2 Evaluation Metric
Existing lane-level accuracy metrics do not reflect global navigation performance. To address this, we propose Association Precision-Recall (Association P-R) for a comprehensive evaluation of the SD-HD map Association quality.
Reachability P-R [22]: This metric evaluates the connectivity of the path between landmarks. A predicted path between locations $\hat{A}$ and $\hat{B}$ is considered a true positive (TP) if its Chamfer distance to any ground-truth path between corresponding $A$-$B$ pairs is under threshold, regardless of direct connectivity between $\hat{A}$-$\hat{B}$.
Association P-R: The Association P-R is introduced with two significant upgrades based on Reachability P-R, as illustrated in Fig. 4:
  1. 1.
    Label Sequence Alignment: Both predicted and ground-truth paths are converted to simplified label sequences representing SD map link traversals.
  2. 2.
    Length-aware Overlap Check: For each aligned label, we calculate the overlap ratio between the predicted and ground-truth path segments. An TP is confirmed when this ratio exceeds the threshold $T$.
Following the mAP conventions [18], we use $T=[0.5:0.05:0.95]$ (10 thresholds) and report mean P-R and F1 scores. To mitigate path-length bias, we separately compute metrics across 15 length intervals $L=[[0,5),[5,10),..,[75,+\infty)]$ before aggregation. Specifically, in OMA-GT, metrics such as A-R and A-F1 become irrelevant because M-R achieves 100% due to the same HD map as the ground truth. The detailed of Association P-R are provided in supplemental material.
5 Method
5.1 Overall Architecture
Figure 5: Overview of Map Association Transformer.
As depicted in Fig. 5, the Map Association Transformer (MAT) is a transformer specifically designed for map association. All inputs are vectorized representations $\mathcal{V}=\{\vec{v}_{1},\vec{v}_{2},\ldots,\vec{v}_{N}\}$, where each vector $\vec{v}_{i}$ is parameterized by two endpoints and direction: $\vec{v}_{i}=[p_{i1}^{x},p_{i1}^{y},p_{i2}^{x},p_{i2}^{y},\theta_{i}]$, with $\theta_{i}=\arctan\left(\frac{p_{i2}^{x}-p_{i1}^{x}}{p_{i2}^{y}-p_{i1}^{y}}\right)$ and $p_{i1},p_{i2}\in\mathbb{R}^{2}$ being the start/end points. The input maps are composed of an SD map, an HD map, and a boundary. The SD map ($\mathcal{G}_{\mathcal{R}}$) comprises road vectors =$\mathcal{R}=\{r_{1},\ldots,r_{m_{r}}\}$, which form a graph with topological edges $\mathcal{E}_{\mathcal{R}}$. Each road $r_{j}$ is transformed into an ordered sequence of vectors $\mathcal{R}_{j}=\overrightarrow{q_{j1}q_{j2}},\overrightarrow{q_{j2}q_{j3}},\ldots$ through its parameterized segments. The HD map ($\mathcal{G}_{\mathcal{L}}$) consists of centerline vectors =$\mathcal{L}=\{l_{1},\ldots,l_{m_{l}}\}$ that represent the centerlines. Each centerline $l_{i}$ is transformed into vectors $\mathcal{L}_{i}=\overrightarrow{p_{i}^{1}p_{i}^{2}},\overrightarrow{p_{i}^{2}p_{i}^{3}},\ldots$ based on consecutive points $p_{i}^{j}$. The boundary ($\mathcal{B}$) includes the boundary vectors =$\mathcal{B}=\{b_{1},\ldots,b_{m_{b}}\}$, converted similarly to the roads: $b_{j}\to\mathcal{B}_{j}=\overrightarrow{h_{j1},h_{j2}},\overrightarrow{h_{j2}h_{j3}},\ldots$. These vectors are processed by the vector embedding module, which maps each 5D vector $\vec{v}_{i}$ to a high-dimensional feature $F_{\vec{v}_{i}}\in\mathbb{R}^{C}$ via a two-layer MLP. The outputs are aggregated into feature matrices: $F_{road}\in\mathbb{R}^{N_{r}\times C}$, $F_{centerline}\in\mathbb{R}^{N_{l}\times C}$, and $F_{boundary}\in\mathbb{R}^{N_{b}\times C}$, where $N_{(\cdot)}$ denotes the total number of vectors (e.g. $N_{r}=\sum_{j=1}^{m_{r}}\text{length}(\mathcal{R}_{j})$).
Subsequently, the vector tokens are input into a transformer network. MAT consists of stacked MAT blocks to extract hierarchical features, each block containing Path-Aware Attention (PA), Spatial Attention (SA), and feed-forward network (FFN). In the association head, each road token requires the pooling to obtain a representative token $\bar{F}_{road}$ for the current road. The association of the centerline with the road is calculated by combining attention between $\bar{F}_{road}$ and $F_{centerline}$, generating the probability distribution of the associations of the SD-HD map. The association probabilities are further refined by a post-processing method to enforce topological constraints. The details of the implementation are provided in supplemental material.
5.2 Path-aware Attention
Attention on path order. Inspired by MapTRv2 [17], we introduce the path-aware attention to iteratively refine the global features of the tokens within the road / centerline graphs. Unlike MapTRv2, which focuses on local path segments, we explicitly model long-range dependencies by constructing paths from root nodes to leaf nodes in the graph. This ensures accurate association across distant vectors, albeit at the cost of redundant computations for overlapping path segments.
In path-aware attention, we first reorder vector tokens to align with their path indices: Given a path token sequence $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$, the vector tokens are rearranged into $\mathcal{V}_{\text{path}}=\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{m}}\}$, where each $v_{i}$ belongs to a sub-path in $\mathcal{P}$. After computing attention over $\mathcal{V}_{\text{path}}$, we apply a scatter-mean operation to project tokens back to the original instance order:
$v_{j}^{\text{inst}}=\frac{1}{|\mathcal{I}_{j}|}\sum_{i\in\mathcal{I}_{j}}v_{i}^{\text{path}},\quad\text{where }\mathcal{I}_{j}=\{i\mid\text{token }v_{i}^{\text{path}}\text{ maps to }v_{j}^{\text{inst}}\}.$ (4)
RoPE in PA. We apply multi-dimensional Rotary Position Embedding (RoPE) [28, 2] to encode positional information in path-aware attention. Specifically, RoPE operates on two dimensions: path-level order, which represents the ordinal position of a vector within its parent path (e.g. $\text{pos}_{\text{path}}\in\mathbb{Z}$), and instance-level order, which denotes the ordinal position of the vector within its local segment (e.g. $\text{pos}_{\text{seg}}\in\mathbb{Z}$). This dual-axis encoding scheme enables the model to disentangle global path structures from local geometric details by explicitly modeling both hierarchical relationships and sequential dependencies among vectors.
5.3 Spatial Attention
Overlook of spatial. Existing transformer models for map processing often neglect explicit spatial coordinates due to their generative nature, making it difficult to associate tokens with physical locations. In contrast, our task benefits from explicit spatial annotations (e.g., GPS coordinates), enabling precise geometric reasoning.
Attention with Vector Serialization. Drawing inspiration from PTv3 [31], we propose a spatial attention mechanism based on vector serialization. Each vector token $\vec{v}_{i}=(p^{1},p^{2})\in\mathbb{R}^{2}$ is encoded in a 3D spatial coordinate $(x,y,r)$, where $x=\lfloor\frac{p^{1}_{x}+p^{2}_{x}}{2g}\rfloor$ and $y=\lfloor\frac{p^{1}_{y}+p^{2}_{y}}{2g}\rfloor$ represent the integer grid coordinates of the vector’s centroid ($g=0.1$m), and $r=\left\lfloor\frac{\theta_{i}}{2\pi/R}\right\rfloor$ denotes the quantized direction angle $\theta_{i}=\arctan\left(\frac{p^{2}_{y}-p^{1}_{y}}{p^{2}_{x}-p^{1}_{x}}\right)$ divided into $R=16$ segments. These coordinates are then assigned to a 1D sequence via a space filling curve $\varphi^{-1}:\mathbb{Z}^{3}\to\mathbb{Z}$, such as the Hilbert curve [11] and the Z curve [24], which preserves spatial location by ordering vectors based on their geometric proximity.
After serialization, tokens are grouped by receptive field size and processed via grouped attention (similar to PTv3). Finally, an inverse serialization operation restores the original token order.
RoPE in SA. In spatial attention, multi-dimensional RoPE encodes absolute spatial positions using $(x,y,r)$, enhancing the model’s ability to focus on geospatial relationships. This complements the relative position bias in standard self-attention.
5.4 Association and Loss Function
Association. The association between the roads and the centerline is calculated through a cross-attention mechanism. For each road $j$, we first aggregate its token features $\{F^{road}_{j1},\ldots,F^{road}_{jN}\}$ into a representative feature $\bar{F}^{road}_{j}=\frac{1}{N}\sum_{n=1}^{N}F^{road}_{jn}$, where $N$ denotes the number of road tokens on roads $r_{j}$ and $F^{road}_{jn}\in\mathbb{R}^{d}$. The association probability $Prob_{ij}$ between centerline $i$ and road $j$ is then calculated as:
$Prob_{ij}=\frac{\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{j}}{\sqrt{d}}\right)}{\sum_{k=1}^{K}\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{k}}{\sqrt{d}}\right)},$ (5)
where $F^{cl}_{i}\in\mathbb{R}^{d}$ is the centerline token feature, $d$ is the feature dimension, and $K$ represents the total number of road. This formulation normalizes the similarity scores in all roads for each centerline $i$, ensuring a valid probability distribution.
Loss Function. We optimize the model using a combination of Cross-Entropy (CE) Loss and Connectionist Temporal Classification (CTC) Loss. The CE Loss monitors the classification of the centerline by maximizing the log-likelihood of ground-truth associations. To enhance topological consistency, we further apply CTC Loss to align centerline token sequences with road structures. Specifically, each centerline token sequence $\{F^{cl}_{i1},\ldots,F^{cl}_{iT}\}$ is treated as a temporal signal, and the CTC loss enforces alignment with the road token sequence $\{F^{road}_{j1},\ldots,F^{road}_{jT}\}$. The total loss is a weighted sum:
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{CE}}+\beta\cdot\mathcal{L}_{\text{CTC}},$ (6)
with hyperparameters $\alpha$ and $\beta$ balancing the two objectives. In practice, $\alpha=1,\beta=0.01$.
5.5 Topology Post-Process
We formalize topological decoding as a structured prediction on all path of the centerline $\mathcal{P}_{j},\ j\in[1,\cdots,K]$. $K$ is the number of total paths. The two-stage decoding process operates as follows:
Token Initialization. For each centerline path $\mathcal{P}_{j}$, we select the initial centerline $T_{\text{max}}$ via:
$T_{\text{max}}=\mathop{\text{argmax}}\limits_{l\in\mathcal{P}_{j}}\mathop{\text{max}}\limits_{r\in\mathcal{R}}P(l,r)$ (7)
where $P(l,r)$ is the probability of association from centerline $l$ to road $r$.
Topological-constraint Beam Search. Based on beem search, topological-constraint beam search makes the following two improvements:
  • Modify the one-way search to implement a bidirectional search starting on $T_{max}$.
  • When generating new predictions, instead of using the approach of taking the maximum value from all roads, we decode under the constraint of connectivity provided in the road network $\mathcal{E}_{r}$, thus ensuring that the connectivity of the road sequence corresponding to the lane path in the decoding result is consistent with the representation of the road network.
Detailed expressions of the topological-constraints Beam Search, including formula descriptions, are included in supplemental material.
6 Experiment
Methods A-F150 A-F175 A-F195 A-P50:95 A-R50:95 A-F150:95 La./ms
KNN [5] 36.8 36.4 32.4 47.9 27.2 34.6 313
HMM [25] 37.6 36.9 32.6 49.6 28.3 36.0 561
MAT(Ours) 43.4 43.0 38.3 56.1 33.9 42.3 74
Table 1: Result on OMA. La. means latency.
      Methods       A-P50       A-P75       A-P95       A-P50:95       La./ms
      KNN [5]       72.2       69.9       59.4       68.6       299
      HMM [25]       73.8       71.5       60.3       70.1       465
      MAT(Ours)       81.6       80.4       69.0       78.7       70
Table 2: Result on OMA-GT. La. means latency.
6.1 Implement detailed
We train our models from the beginning for a total of 50 epochs utilizing the AdamW optimizer. A cosine-decay learning rate scheduler is employed, incorporating a linear warm-up phase lasting for one epoch. The starting learning rate, weight decay, and batch size are set at 0.0001, 0.05, and 128, respectively, using a NVIDIA A6000 GPU. The latency of MAT operates on an NVIDIA A6000 with pytorch, while KNN and HMM function on an Intel(R) Xeon(R) Platinum 8369B. The KNN and HMM code is developed using Python and Numpy without optimization. In practical applications, highly optimized C++ code could provide better time efficiency.
6.2 Result
[…]
6.3 Ablation Study
The ablation study experiment is conducted within the OMA-GT dataset, with latency measured using a NVIDIA A6000.
Structure Input Post. Loss Road Pooling Metric
No. PA SA ROPE Boundary Post. CE CTC Avg. Max A-P50:95 La./ms
1 74.1 61
2 62.1 77
3 77.8 64
4 77.9 68
5 78.5 69
6 78.4 70
7 67.7 70
8 78.5 70
9 78.7 70
Table 3: Ablation study of structure, input, loss function and road pooling method. Post. means post process. Avg. means average pooling. La. means latency.
Patch Size 64 256 1024 2048 $\infty$
A-P50:95 77.8 78.4 78.7 78.5 78.4
Latency/ms 77 68 70 72 72
Table 4: Ablation study of patch size of spatial attention.
Group Method N.G Category Path
A-P50:95 78.0 78.2 78.7
Latency/ms 75 73 70
Table 5: Ablation study of group method of path attention. N.G. means non-grouping.
[…]
Visualization and Failed cases. The visualizations and failed cases have been included in the supplemental material. Our approach, as demonstrated by the visualization results, facilitates a stronger long-range correlation on OMA-GT and ensures dependable SD-HD alignment on OMA.
7 Conclusion
This work introduces OMA, the first dataset for online map association with annotated correspondences with a path-based association Precision-Recall metric aligned with navigation requirements. Furthermore, we introduce MAT, a transformer framework with dual attention mechanisms for map association. MAT achieves 32.3% A-F150:95 / 78.7% A-P50:95 on OMA / OMA-GT with 74 / 70 ms latency, outperforming HMM by 5.5% / 8.6%. The limitation of OMA is its exclusion of dynamic components, such as traffic lights, that affect real-world navigation. In future studies, we plan to incorporate these environmental data into the dataset and assess how it impacts the accuracy of the association.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62433003) and the National Natural Science Foundation of China (62476017).
Appendix A Dataset Analysis
Revised Analysis. As detailed in Section 4.1, the dataset is partitioned into OMA-GT (ground-truth HD maps) and OMA (predicted HD maps), with statistics summarized in Tab. 6. OMA-GT comprises 26,111 training scenarios and 5,613 validation scenarios, totaling 31,724 samples, while OMA contains only 5,573 test scenarios due to the exclusion of low-quality predictions. Both data sets share identical spatial coverage, with HD maps covering (±$(\pm 15m,\pm 30m)$ and SD maps extending to (±$(\pm 75m,\pm 75m)$. Notably, the SD map’s road density in OMA-GT decreases from 15.1 roads/scene during training to 10.8 in validation/test splits, suggesting potential domain shifts between training and evaluation environments.
Quantitative discrepancies between OMA and OMA-GT reveal systemic geometric and topological inconsistencies in predicted maps. OMA predicts an average of 310 lanes/scene, more than four times that of OMA-GT (78.6), with significantly shorter mean lane lengths (2.32 m vs 3.16 m in OMA-GT), indicating both oversegmentation and false positives. This fragmentation is further amplified by OMA’s prediction of 322.06 lane paths/scene (vs. 7.69 in OMA-GT), where ground-truth lanes are frequently split into disconnected fragments. Boundaries exhibit similar degradation: OMA detects 9.65 boundaries/scene (vs 3.40 in OMA-GT) with reduced mean lengths (32.17m vs 44.23m), reflecting fragmented boundary detection. Meanwhile, OMA-GT validation data show slight degradation compared to training splits (e.g. 75.8 vs. 81.3 lanes/scene), highlighting inherent variability in real-world map quality.
Connectivity metrics expose deeper structural errors in OMA predictions. The average lane connectivity in OMA reaches 2.9, substantially higher than OMA-GT’s 2.0/2.1, revealing widespread mislinking of spatially disjoint lanes. Similarly, the OMA-GT validation data show reduced road connectivity (1.8 vs. 2.0 in training), suggesting a domain bias toward simpler topologies in training scenarios. Semantic associations between roads and lanes also degrade significantly. OMA-GT training roads are associated with 1,547 lanes on average, collapsing to 945 in validation splits, which implies degradation of the hierarchical structure in complex scenarios.
These discrepancies have critical implications for benchmarking perception systems. The severe overprediction and fragmentation in OMA highlight the need for metrics penalizing false positives and disconnected paths (e.g., path-length-weighted scores). Furthermore, the mismatch between OMA-GT’s training and validation/test distributions (e.g., road count/length differences) necessitates domain adaptation strategies to ensure generalization. Finally, the collapse of semantic hierarchies in validation data suggests that end-to-end models may struggle to learn robust associations between roads and their constituent elements without explicit structural constraints. Together, these findings underscore the importance of a connectivity-aware association method to avoid overestimating performance on fragmented or mislinked predictions.
Pon split. Both OMA-GT and OMA apply the pon split [27] of the nuScenes dataset [4], ensuring that there is no leakage between the training and validation datasets. For consistent lane prediction segmentation with OMA, we re-trained MapTRv2 [17] using the nuScenes dataset with the pon split. However, this resulted in a significant drop in the quality of the centerline network prediction by MapTRv2 when using the pon split, which adversely affects OMA’s current metrics. Drawing inspiration from the private protocol in MOT17/MOT20 [7], we suggest that future research evaluates the OMA dataset with an enhanced centerline prediction network, without relying on MapTRv2 as a baseline.
Appendix B Visualization
This section presents the visualization of our model, featuring path-aware attention (PA) and spatial attention (SA) alongside the model’s results. Furthermore, we examine the failed case with an analysis of our model.
Figure 6: Visualization of attention map of Path-aware attention and spatial attention. SA means Spatial Attention. PA mean Path-aware Attention. The red triangle represents the token corresponding to the current attention map.
B.1 Attention Map
The upper portion of Fig.6 presents visualizations of Spatial Attention (SA) maps at different stages of the model. As revealed by the analysis, SA provides extensive receptive fields in the early stages, enabling tokens to capture global contextual information. Specifically, during Stage 1 and Stage 2, the SA attention distributions exhibit highly dispersed patterns, allowing each query token to uniformly attend to global regions across the input space. In later stages, the functional role of SA transitions to facilitating cross-category token interactions. For example, in Stages 3-5, distinct attention patterns emerge where tokens primarily interact with their semantically corresponding road elements. Notably, this interaction is not strictly confined to the road tokens directly associated with the centerline token - significant attention weights also develop between the centerline token and adjacent road segments. As exemplified in Stage 4, the tokens establish prominent attention links with multiple road tokens along the same path. We posit that this expanded interaction mechanism constitutes a critical component for precise centerline localization. The propagation of attention observed in later stages effectively enables the maintenance of geometric coherence between spatially distributed road elements while preserving discriminative semantic information through long-range dependencies.
In contrast, the lower part of Fig.6 visualizes the Path-aware Attention (PA) maps at different stages of the model. The visualization reveals that PA primarily focuses on neighboring tokens adjacent to the target path tokens, effectively serving as a local information extractor. Experimental results demonstrate that this localized information extraction capability plays a pivotal role in the model performance, exhibiting a marked contrast with the global perception mechanism of SA. We posit that SA specializes in capturing global contextual patterns while PA emphasizes localized feature extraction. This dual-attention paradigm establishes a synergistic interplay between global and local perception, achieving an optimal balance between comprehensive understanding and fine-grained detail processing, thereby substantially enhancing the model’s overall effectiveness.
Figure 7: Visualization of result in OMA-GT.
Figure 8: Visualization of result in OMA.
B.2 Result Compare
Fig.7 illustrates comparisons of model ground truth on OMA-GT, KNN, HMM, MAT and MAT w/ postprocess. The top rows of (d) and (e) exhibit our post-processing module’s effectiveness. MAT predictions initially display incorrect topological connections where roads are mistakenly linked (marked with red circles). Our postprocessing, which utilizes topology-aware beem searching, rectifies this by eliminating non-sequential transitions and reconstructing precise topological paths. The second row demonstrates MAT’s superior handling of complex topologies. Although HMM targets sequential path associations, its single path paradigm often underperforms in complex topologies with intersections. In contrast, our model uses spatial attention to grasp global information and cross-path associations, facilitating adaptive learning of complex topological patterns for accurate connectivity inference. The third row showcases our model’s improved ability to localize associations. Using path-aware attention, the model emphasizes detailed extraction of local features along paths. This targeted local perception ensures precise associations at challenging points, such as junctions, where HMM is typically short due to limited contextual understanding.
Fig. 8 illustrates a comparison of ground truth results for OMA-GT, KNN, HMM, MAT, and MAT with post-processing. Significantly, the visualization demonstrates that our model excels in map association in noisy scenarios with inaccurate centerline predictions, surpassing KNN and HMM by integrating the complementary benefits of global association (SA) and local detail refinement (PA).
Figure 9: Visualization of failed cases.
B.3 Failed Cases
Fig. 9 illustrates the failed cases of the MAT. Our study reveals that the key challenge is the localization errors associated with spatial misalignment between the predicted paths and the actual labels. This discrepancy significantly affects the accuracy of the association, particularly at critical junctures where complex path interactions create ambiguous topological patterns. Although all baseline methods exhibit substantial association errors under these difficult conditions, our model achieves notable error reduction due to its dual-attention framework. However, discrepancies between our predictions and the ground truth remain, indicating potential for further enhancement. We propose that improving the path-aware attention (PA) mechanism by incorporating local operators such as convolutional kernels could be advantageous. This hybrid approach would preserve model efficiency while allowing for more precise spatial-temporal feature extraction at path intersections, thus improving local association accuracy without compromising inference speed.
Appendix C Model Families
Methods A-P50 A-P75 A-P95 A-P50:95 La./ms
KNN [5] 72.2 69.9 59.4 68.6 299
HMM [25] 73.8 71.5 60.3 70.1 465
MAT-T 81.2 80.0 68.1 78.2 34
MAT-S 81.4 80.2 68.3 78.3 56
MAT-M 81.5 80.3 68.9 78.6 63
MAT-L (Main Manuscript) 81.6 80.4 69.0 78.7 70
Table 7: Result on OMA-GT. La. means latency.
We develop a family of models derived from those detailed in the main manuscript. By adjusting the number of layers and channels, these models vary in size, memory usage, and inference speed, allowing them to suit various end-side environments. Detailed parameters for each model can be found in Section F of the Supplementary Material.
The experimental results in Tab. 7 demonstrate the effectiveness of our MAT architecture in the accuracy and efficiency dimensions. As the model scales from MAT-T to MAT-L, we observe a consistent improvement in precision metrics (A-P50: 81.2$\rightarrow$81.6, A-P75: 80.0$\rightarrow$80.4, A-P95: 68.1$\rightarrow$69.0) with increasing parameter counts, although at the cost of reduced inference speed (34$\rightarrow$70ms). In particular, even the smallest variant MAT-T outperforms HMM by achieving a 9.6% higher A-P50:95 accuracy while reducing latency by 92.6%, establishing a new Pareto frontier for trajectory association. The minimal performance degradation from A-P50 to A-P95 (-12.6% for MAT-L vs -13.5% for HMM) further highlights the robustness of our dual attention mechanism in preserving topological constraints under stringent matching criteria. The family of MAT allows users to select optimal configurations based on hardware constraints, with potential future extensions toward dynamic computation allocation that activates larger models only during complex topological transitions. These findings quantitatively confirm that explicit modeling of both global topology (SA) and local geometry (PA) through differentiable attention mechanisms can surpass traditional probabilistic approaches while satisfying practical deployment requirements.
Appendix D Metric Details
In the main article, we present a narrative explanation of the Association P-R accompanied by a schematic diagram. To elucidate the calculation of Association P-R more thoroughly, we include the pseudo-code for computing Association P-R, as depicted in Alg. 1.
Algorithm 1 Evaluate Association P-R
1:function EvalMetric(pred_centerline, gt_centerline, threshold, acc_list)
2:     Input: pred_centerline, gt_centerline, threshold, acc_list
3:     Step 1: Point Matching
4:     ExtractPoints(pred_centerline)
5:     ExtractPoints(gt_centerline)
6:     PointMatch(pred_sample, gt_sample_point, threshold)
7:     Step 2: Path Matching
8:     InitializeCounters(TP, FP, FN, acc_list)
9:     for all point pairs $(i,j)$ in matched points do
10:         Find pred_path and gt_path between points $i,j$
11:         PathMatch(pred_path, gt_path, threshold)
12:         if paths match then
13:              Check sequence consistency and accuracy
14:              for all $acc\in acc\_list$ do
15:                  Update TP/FP based on accuracy vs $acc$
16:              end for
17:         end if
18:     end for
19:     Step 3: Count Unmatched Paths
20:     for all unmatched gt path do
21:         for all $acc\in acc\_list$ do
22:              $FN[acc][k]\leftarrow FN[acc][k]+1$
23:         end for
24:     end for
25:     Step 4: Calculate Precision and Recall
26:     Initialize: $Precision\leftarrow\{\},Recall\leftarrow\{\}$
27:     for all $acc\in acc\_list$ do
28:         $denominator\_p\leftarrow TP[acc]+FP[acc]$
29:         $denominator\_r\leftarrow TP[acc]+FN[acc]$
30:         $Precision[acc]\leftarrow\begin{cases}TP[acc]/denominator\_p&\text{if }denominator\_p>0\\ 0&\text{otherwise}\end{cases}$
31:         $Recall[acc]\leftarrow\begin{cases}TP[acc]/denominator\_r&\text{if }denominator\_r>0\\ 0&\text{otherwise}\end{cases}$
32:     end for
33:     Return: TP, FP, FN, Precision, Recall
34:end function
[…]
$\begin{split}&\text{A-P}^{50:95}=\sum_{th\in T}\text{A-P}^{th},\quad\text{A-R}^{50:95}=\sum_{th\in T}\text{A-R}^{th}\\ &\text{A-F1}^{th}=\frac{2\text{A-P}^{th}\cdot\text{A-R}^{th}}{\text{A-P}^{th}+\text{A-R}^{th}},\quad\text{A-F1}^{50:95}=\frac{2\text{A-P}^{50:95}\cdot\text{A-R}^{50:95}}{\text{A-P}^{50:95}+\text{A-R}^{50:95}}\end{split}$ (8)
where $th$ are the thresholds in association P-R as $[0.5:0.05:0.95]$ (10 thresholds).
Appendix E Model Details
In this section, we delve deeper into the technical aspects of the model, including Path-aware attention, spatial attention, and the model’s post-processing.
Figure 10: Overview of Path-aware attention and spatial attention.
E.1 Path-aware attention
The particular design of Path-aware attention (PA) can be seen in Fig 10 (a). The fundamental framework of PA is made up of four components: computing order and its inverse, reorganizing tokens, calculating attention, and inverting tokens.
Path-aware attention uses paths to determine the sequence of tokens. Initially, we define the network of roads or centerlines and then identify all complete paths from a starting point (with no incoming connections) to an endpoint (with no outgoing connections). We concatenate these paths to generate a path-based token sequence. During the reordering of path-aware attention, a token may appear across various paths simultaneously, requiring us to duplicate the token. Then, we compute attention by segregating tokens based on their paths, ensuring that interactions occur only among tokens within the same path. After attention calculation, the tokens are reversed to match the original input sequence. If multiple tokens exist within the path of a single original token, they are averaged.
E.2 Spatial attention
Fig. 10 (b) provides a detailed description of the architecture of the SA model. Similarly to pa, the fundamental structure of SA includes four main components: computing sorting and its reverse, rearranging tokens according to the sorted order, calculating attention, and then reverse sorting the tokens again.
In order of SA, we apply a space filling curve $\varphi^{-1}:\mathbb{Z}^{3}\to\mathbb{Z}$ to serialize the vectors in a 1D sequence, preserving spatial locality. Four curves are used: Z-order, Transposed-Z, Hilbert, and Transposed-Hilbert. To avoid bias toward specific curve types, we randomly select one curve per training iteration. In addition, similar to PA, if there are multiple tokens that belong to a single token after the coordinate calculation, we will average the multiple tokens in the sort and copy that token to all the corresponding tokens in the reverse sort.
E.3 Post Process
In the main manuscript, we offer a narrative explanation of the post-processing. To enhance clarity, we also present a mathematical formulation of the post-processing details. Let $\mathcal{R}$ denote the road as the vocabulary in the traditional beem search with size $|\mathcal{R}|$, and $k$ represent the beam width. At each step $t$, the algorithm maintains a set $B_{t}$ of candidates path $k$, each associated with a score $s(h)$ defined as the sum of logarithmic conditional probabilities. The search begins by selecting the token $w^{*}$ with the maximum initial probability $P(w|x)$ given as input $x$, forming the singleton initial set:
$B_{0}=\text{Top}_{1}\left(\mathcal{R},\ \log P(w|x)\right),$ (9)
which simplifies to:
$B_{0}=\left\{[w^{*}]\right\},\ \text{where}\ \log P(w^{*}|x)=\max_{w\in\mathcal{R}}\log P(w|x).$ (10)
This initialization bypasses conventional fixed start tokens and prioritizes high-probability seeds.
At iteration $t\geq 1$, each sequence $h\in B_{t-1}$ generates $2|\mathcal{R}|$ candidates by appending a token $w\in\mathcal{R}$ to either the left ($w\cdot h$) or right ($h\cdot w$) of $h$, forming the expanded candidate set:
$\mathcal{C}_{t}=\left\{w\cdot h\mid h\in B_{t-1},\ w\in\mathcal{R}\right\}\cup\left\{h\cdot w\mid h\in B_{t-1},\ w\in\mathcal{R}\right\}.$ (11)
h extension updates the sequence score using direction-specific conditional probabilities:
$s(h^{\prime})=\begin{cases}s(h)+\log P(w|h,\text{left},x),&\text{if }h^{\prime}=w\cdot h\\ s(h)+\log P(w|h,\text{right},x),&\text{if }h^{\prime}=h\cdot w\end{cases}.$ (12)
The top-$k$ candidates from $\mathcal{C}_{t}$ are retained to form $B_{t}$:
$B_{t}=\text{Top}_{k}\left(\mathcal{C}_{t},\ s(\cdot)\right),$ (13)
i.e.,
$B_{t}=\left\{h^{\prime}_{1},h^{\prime}_{2},\dots,h^{\prime}_{k}\right\},\ \text{with}\ s(h^{\prime}_{1})\geq s(h^{\prime}_{2})\geq\dots\geq s(h^{\prime}_{k}).$ (14)
The process terminates at a predefined maximum length $T$ or when all sequences emit an end-of-sequence token, with the final output $\hat{h}$ selected as:
$\hat{h}=\arg\max_{h\in\bigcup_{t=0}^{T}B_{t}}s(h).$ (15)
Appendix F Implement Details
F.1 Model Settings
Tab. 8 summarizes the architectural configurations of our proposed MAT variants (MAT-T, MAT-S, MAT-M, MAT-L). All variants adopt identical channel dimensions ([96,192,384,768,1536]), attention head counts ([4,4,8,8,8]), spatial curve orders (["z", "z-trans", "hilbert", "hilbert-trans"]) and hybrid attention mechanisms combining spatial attention (SA) and path-aware attention (PA). In particular, parameters such as patch sizes (1024), MLP ratios (matching spatial curve orders), and stochastic depth rates (0.3) are uniformly inherited across architectures, reflecting ablation study results that optimized these values for balanced accuracy-latency trade-offs. A distinctive design choice lies in the shuffling strategy, where MAT-T/S/M/L progressively refine the shuffle order to enhance token mixing in spatial attention, aligning with their increasing computational budgets. This structured configuration hierarchy enables systematic evaluation of model capacity versus efficiency, as validated by the ascending La / ms metrics (34 $\rightarrow$ 70) corresponding to deeper transformer layers.
Parameter MAT-T MAT-S MAT-M MAT-L
Blocks [2, 2, 2, 2, 2] [4, 4, 4, 4, 4] [4, 4, 4, 8, 4] [4, 4, 4, 12, 4]
Attention Head [4, 4, 8, 8, 8]
MLP Ratio [4, 4, 4, 4, 4]
Drop Path [0.3, 0.3, 0.3, 0.3, 0.3]
Channels [96, 192, 384, 768, 1536]
Path Size [1024, 1024, 1024, 1024, 1024]
Attention Order ["Spatial Attention", "Path-aware Attention"]
Spatial Curve ["z", "z-trans", "hilbert", "hilbert-trans"]
Shuffle [Shuffle Order, Shuffle Order, Shuffle Order, Shuffle Order, Shuffle Order]
Latency/ms 34 56 63 70
Table 8: Model settings.
Training Configuration
optimizer AdamW batch size 128
scheduler Cosine weight decay 5e-3
learning rate 1e-5 epochs 50
block lr scaler 1e-5 warmup epochs 2
criteria CrossEntropy, CTC Loss
Data Augmentation
random rotate axis: z, angle: [-1, 1], p: 0.5 random scale scale: [0.9, 1.1]
random flip p: 0.5 random jitter sigma: 0.005, clip: 0.02
grid sampling grid size: [0.1, 0.1, $\pi/16$]
Table 9: Train Configuration and Data augmentations.
F.2 Train Configuration
The details of the implementation are summarized in Tab. 9. Our training protocol employs the AdamW optimizer with a base learning rate of $2\times 10^{-3}$ and a cosine learning rate decay, operating in mini-batches of size 12. The weight decay regularization is set to $5\times 10^{-3}$, while the block-wise learning rate scaling ($10^{-1}$ factor) is applied to stabilize the propagation of the gradient across the transformer layers. The training process spans 50 epochs with a 2-epoch warm-up phase for learning rate initialization. Model optimization combines CrossEntropy loss for classification tasks and CTC loss for sequence alignment objectives.
F.3 Data Augmentations
For data augmentation as shown in Tab. 9, we implement a series of randomized transformations that include axis-aligned rotation around the z-axis with ±$\pm 1^{\circ}$ angular variation at a 50% application probability, isotropic scaling within the range $[0.9,1.1]$, random flipping with equal probability 50%, point cloud jittering characterized by $\sigma=0.005$ and a clip limit of 0.02, and grid sampling with spatial discretization parameters set to $[0.1,0.1,\pi/16]$. These enhancement strategies were systematically validated through ablation studies to optimize the balance between model accuracy and computational efficiency while ensuring robustness to input variations.